Adult tickets for a play cost $16 and child tickets cost $3. If there were 23 people at a performance and the theater collected $199 from ticket sales, how many adults and how many children attended the play?
Found 2 solutions by checkley75, heidi:
Answer by checkley75(3666)

16A+3C=199
A+C=23 OR A=23-C
NOW SUBSTITUTE (23-C) FOR A & SOLVE FOR C
16(23-C)+3C=199
368-16C+3C=199
-13C=199-368
-13C=-169
C=-169/-13
C=13 IS THE NUMBER OF CHILDREN'S TICKETS SOLD.
23-13=10 ADULT TICKETS SOLD.
PROOF
16*10+13*3=199
160+39=199
199=199

Answer by heidi(17)

there are 7 adults and 29 children


cause: 16*7 = 112 , $199-$112= 87 , 87/3= 29